Please Help!
Determine the mean, median, mode, and midrange of the set of data. Round your answer to the nearest tenth.
a) 7,5,8,8,8,10,12
b) 1,1,1,1,4,4,4,4,6,8,10,12,15,21
The mean represents the average or the sum of the data divided by the number of pieces of data
The median is the value in the middle of a set of ranked data
The mode is the piece of date that occurs most frequently
The midrange is the value half way between the lowest and highest
values in a set of data. It is found by adding the loweast and highest values and dividing the sum by 2.
This is all I have been able to figure out about this problem-please help me piece it together!
Thank you in advance

Answer by stanbon(75887)

Determine the mean, median, mode, and midrange of the set of data. Round your answer to the nearest tenth.
==============================
a) 7,5,8,8,8,10,12
The mean represents the average or the sum of the data divided by the number of pieces of data : mean = 58/7 = 8.2857
--------
The median is the value in the middle of a set of ranked data
median = 8
-------------------
The mode is the piece of date that occurs most frequently
mode = 8
-----------------
The midrange is the value half way between the lowest and highest
values in a set of data. It is found by adding the loweast and highest values and dividing the sum by 2.
midrange = (5+12)/2 = 8.5
====================================
==================================
b) 1,1,1,1,4,4,4,4,6,8,10,12,15,21
The mean represents the average or the sum of the data divided by the number of pieces of data
mean = 92/14 = 6.57...
-----------
The median is the value in the middle of a set of ranked data
median = 4
-------------
The mode is the piece of date that occurs most frequently
You have two modes: 1 and 4 ; the data is said to be bimodal
-----------------
The midrange is the value half way between the lowest and highest
values in a set of data. It is found by adding the loweast and highest values and dividing the sum by 2.
midrante = (1+21)/2 = 11
=============
Cheers,
Stan H.